x 2 y 2 dx+1/xdy=0

x2y2dx+1/xdy=0

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Jithi

x2y2dx+1/xdy=0

 x2y2dx+=-1/x dy

x3 dx = -dy/y2

x3 dx =-1/y2 dy

x4/4  = 1/y  +c

 

Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin